Tucson Bank offers to lend you $50,000 at a nominal rate of 12%, compounded monthly. The loan (principal plus interest) must be repaid at the end of the year. Phoenix Bank also offers to lend you the $50,000, but it will charge an annual rate of 10.8%, with no interest due until the end of the year. How much higher or lower is the effective annual rate charged by Tucson versus the rate charged by Phoenix?

Solve without Excel. Show equations and steps.
Answer #1

Tucson Bank EAR=(1+r/m)^m-1=(1+12%/12)^12-1=12.68%

Phoenix Bank EAR=10.8%

Tucson rate is higher by 12.68%-10.80%=1.88%